OpenStudy (anonymous):

HELP!!!!! A turntable must spin at 33.7 rpm (3.53 rad/s) to play an old-fashioned vinyl record. How much torque must the motor deliver if the turntable is to reach its final angular speed in 3.0 revolutions, starting from rest? The turntable is a uniform disk of diameter 30.5 cm and mass 0.25 kg.

OpenStudy (anonymous):

Wf=W0+(Angular acceleration)t.....(1) theta=W0t+1/2(Angular acceleration)t^2...(11)....here Wo=0,Wf=3.53rad/sec,theta=3*2pi, solve (1) and (11) and find Angular acceleration.. Fore torque ,apply this equation... torque=moment of inertia*Angular acceleration.... here moment of inertia=1/2mr^2..... m=.25 kg, and r=(30.5*10^-2)/2m..find torque.